# A projection method for particle resampling

### v5

@@ Abstract
This work is motivated by the Vlasov-Maxwell-Landau model of magnetized plasmas with up to six dimensions, $3X$ in physical space and $3V$ in velocity space, and is developed in the context of a $1X$ + $1V$ Vlasov-Poisson model of Landau damping with logically regular particle and continuum phase space grids. Stable long time dynamics are demonstrated up to $T=500$ and reproducibility artifacts and data with stable dynamics up to $T=1000$ is publicly available.
@@ 4.2 Particle grids and finite element order
5 Finite element version of the direct remap method
Colella et al. developed a method that can be viewed as a version of our algorithm in a finite volume context. In this algorithm continuum space points define the new particle points directly allowing for a trivial map from the grid to particles [43, 32, 33]. That is, $M_{p}=I$ in (3) and the pseudoinverse vanishes. Cubic splines are used for particle deposition on Cartesian grids, demonstrate high-order convergence [32], and use phase-space adaptive mesh refinement (AMR) grids [33]. This method should conserve energy because the cubic spines can represent a second order polynomial exactly and the direct remap obviously conserves all moments.
The test harness supports a direct map method in a vertex-centered, finite element context. Linear $Q1$ elements are used in the velocity and space dimension on Cartesian grids. Particles are placed at vertices and the domain is doubly periodic (to make $M_{p}$ square), which is immaterial given that there should be negligible density at the velocity boundary. The salient feature of this construct is that the particle mass matrix is the identity, $\overline{M}_{p}=I$, and the pseudoinverse vanishes. These experiments are intended to observe the qualitative effect of high-order grid to particle mapping. Quantitatively, the direct map approach conserve energy to about four digits with $Q1$ in velocity space (see the src/A.01 directory in the repository). These $Q1$ elements in velocity space with two particles per cell should, in theory conserve momentum only.
Figure 2: Particle electric field amplitude linear Landau damping, $A=0.01$, no resampling (blue and noisy), a direct remapping finite element version of Myers et al. (magenta), and the projection method (red)
Direct map resampling is tested with a linear Landau damping, $A=0.01$, problem with a $128\times 256$ particle grid and $V_{max}=6.0$ and results in Figure 2 agree well with amplitude reported in [32] (Figure 3.1). $Q_{1}-Q_{1}$ is used for the projection method with one particle per cell, or vertex, to mimic the direct map method, and the projection method uses a $2\times 2$ particle grid per cell and half as many cells in each dimension to maintain the same number of particles for all three tests and to ensure stability of the pseudo-inverse. This data shows the efficacy of resampling in that it suppresses the noise observed without resampling for the $A=0.01$ case that was also observed by Colella et al. and some increase in stability is observed with the pseudoinverse technique.
@@ Acknowledgments
This work was supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences and Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) Partnership programs under Contract No. DE-AC02-05CH11231 at Lawrence Berkeley National Laboratory, the Office of Naval Research and by an appointment to the NRC Research Associateship Program at the U.S. Naval Research Laboratory, administered by the Fellowships Office of the National Academies of Sciences, Engineering, and Medicine. DSF and MGK were partially supported by NSF CSSI grant 1931524.

### v6

@@ Abstract
This work is motivated by the Vlasov-Maxwell-Landau model of magnetized plasmas with up to six dimensions, $3X$ in physical space and $3V$ in velocity space, and is developed in the context of a $1X$ + $1V$ Vlasov-Poisson model of Landau damping with logically regular particle and continuum phase space grids. Stable long time dynamics are demonstrated up to $T=500$ and reproducibility artifacts and data with stable dynamics up to $T=1000$ are publicly available.
@@ 4.2 Particle grids and finite element order
5 Direct remap method with low-order finite elements
Colella et al. developed a method that can be viewed as a version of our algorithm in a finite volume context. In this algorithm, the continuum grid data after particle deposition, is mapped directly to new particles [43, 32, 33]. Cubic splines are used for particle deposition on Cartesian grids. They demonstrate high-order convergence [32], and use phase-space adaptive mesh refinement (AMR) grids [33]. This method should conserve energy because the cubic spines can represent a second order polynomial exactly and the direct remap obviously conserves all moments.
The test harness supports a direct map method in a vertex-centered, finite element context with $Q1$ elements. The grid vertices define the new particles and the domain is doubly periodic (to make $M_{p}$ square), which is immaterial given that density is negligible at the velocity boundary. The salient feature of this construct is that the particle mass matrix is the identity, $\overline{M}_{p}=I$, and the pseudoinverse vanishes. These experiments are intended to observe the qualitative effect of, in some sense, high-order particle mapping.
Figure 2: Field amplitude with linear Landau damping, $A=0.01$: no resampling (blue and noisy), a direct remapping finite element version of Myers et al. (magenta), and the projection method (red)
Figure 2 show the results of this direct map algorithm with the psuedoinverse project on a linear Landau damping, $A=0.01$, problem with a $128\times 256$ particle grid and $V_{max}=6.0$ and results in agree well with amplitude reported in [32] (Figure 3.1). $Q1$ elements are used for all spaces. This direct map method requires one particle per cell and the projection method requires two particles per cell in each dimension to ensure stability of the pseudoinverse. To maintain the same number of particles for all three tests, the projection method uses a $2\times 2$ particle grid per cell and half as many cells in each dimension. First, this data shows again, with $A=0.01$ the efficacy of resampling in that it suppresses the noise observed without resampling. There appears to be some increase in stability with the pseudoinverse technique, however the coarser FE grid may provide smoothing.
@@ Acknowledgments
This work was supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences and Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) Partnership programs and the FASTMath Institute under Contract No. DE-AC02-05CH11231 at Lawrence Berkeley National Laboratory, the Office of Naval Research and by an appointment to the NRC Research Associateship Program at the U.S. Naval Research Laboratory, administered by the Fellowships Office of the National Academies of Sciences, Engineering, and Medicine. DSF and MGK were partially supported by NSF CSSI grant 1931524.